(1 point) Consider the linear system = [3] -3 -2 5 3 y. -3-1 a. Find the eigenvalues and eigenvectors for the coefficient matI am not sure about the eigen vectors or the eigen values would like confirmation and the solutions for part B as well, Thank you.
Answer #1

If you have any doubts in the solution please ask me in comments ....here i used basic concept of ode solution

-3 -2 gio) S9 sen ; g = 3] 7 5 -3 -2 where y = ve Dy;] A = 92 5 3 ana 2,1t) Let Solution of System 0 is ü (t) = Cie N & Ge wsolue IUP Now to - - 34, - 242 y, Y 4,(0) = -10 92 (0)=15 sy, + 342 j wrote in system form A is By previous part solution-10 cost » ū(t) = 15 cost - 5 sint. -10 cost, > 9,(t) = yz(t) = ] Answer. 15 wst - 5 sint.